# MedLA: A Logic-Driven Multi-Agent Framework for Complex Medical Reasoning with Large Language Models

### v1

@@ 1. Introduction
LLM has demonstrated significant advantages in the field of medical reasoning, and its deep learning architecture can efficiently extract knowledge from massive amounts of literature and clinical cases to provide intelligent assistance in diagnostic decision-making, and is expected to significantly improve the accessibility and popularization of medical knowledge(Chang et al. 2024; Kasneci et al. 2023). Answering complex medical questions with large language models (LLMs) remains difficult. A practical system must integrate domain knowledge (Liu et al. 2024a), patient information (Yang et al. 2022), and explicit logical reasoning (Goh et al. 2024). Recent general-purpose and domain-adapted models achieve strong open-benchmark scores (Kim, Wang et al. 2024; Tang et al. 2024). In clinical use, however, they may hallucinate drug dosages, misapply guidelines, or draw invalid causal links, reducing diagnostic reliability (Liu, Li et al. 2023).
LLM-based medical reasoning follows two main paradigms. (a) Knowledge fine-tuning (Singhal et al. 2023; Wang et al. 2025) retrains models on large medical corpora, improving accuracy at the cost of data, compute, and deployment agility. (b) Reasoning stimulation (Liévin et al. 2024) has been tried using multi-agent role-playing to accomplish tasks through discussion and cooperation, which is considered a flexible and low-cost solution (Moor, Huang et al. 2023; Kim et al. 2024). However, we found that most current multi-agent systems only engage in positional discussions based on their rulings and cannot argue about logical details in depth. Current frameworks are not able to effectively localize logic/rule conflicts, thus be difficult to improve the performance.
Inspired by the ‘major premise-minor premise-conclusion’ paradigm of the classical syllogism, we use the syllogism as a minimal reasoning unit  (Khemlani and Johnson-Laird 2012; Jiang and Yang 2023). Each triad consists of a generalized medical law (major premise), a patient-specific fact (minor premise), and the corresponding conclusion. By concatenating or parallelizing multiple triads, we construct an inference tree whose leaf nodes store empirical observations or domain rules, internal nodes hold intermediate inferences, and the root node gives the final clinical decision  (Howson 2005). Such inference trees offer two major advantages: (i) Traceability - each conclusion can be traced back to its supporting premises; (ii) Comparability - each intelligence can align the reasoning tree (logical tree) to pinpoint conflicts or omissions at the premise level. Embedding this explicit structure into a multi-agent framework provides an easily auditable logical template for systematic cross-intelligence error correction and consistency verification.
@@
The experimental results show that our MedLA outperforms existing MedDDx benchmarks (Su et al. 2025a), medical QA benchmarks (Xiong et al. 2024), and medical reasoning benchmarks (Zuo et al. 2025) by a large margin on both open-source and commercial LLM (Fig. 1(b)). Contributions. (a) We introduce MedLA, the first multi-agent framework for medical reasoning that represents each agent’s thought process as an explicit logical tree. This design enables fine-grained traceability of inferences and systematic detection of premise-level conflicts. (b) We develop a multi-round, graph-guided discussion mechanism in which agents iteratively compare and revise their logical trees, leading to robust cross-agent error correction and convergence to high-confidence, self-consistent reasoning structures. (c) We perform comprehensive evaluations on both differential diagnosis (MedDDx) and standard medical QA benchmarks, demonstrating that MedLA outperforms static role-based multi-agent systems and single LLM baselines.
2. Related Work
Multi-Agent Systems for Collaborative Problem Solving. Multi-agent systems (MAS) have emerged as effective tools for complex, collaborative tasks, particularly in medicine (Li et al. 2024). Systems like MedAgents adopt fixed specialist roles (e.g., radiologist, pathologist) to simulate clinical collaboration for zero-shot QA (Tang et al. 2024), while MDAgents employs hierarchical coordination to enhance flexibility and diagnostic accuracy (Kim, Wang et al. 2024). Others, such as GopherCite (Rae et al. 2021), integrate retrieval-augmented generation to infuse domain knowledge. However, most rely on static role structures, limiting adaptability to novel or evolving scenarios (Guo et al. 2024).
Logic-Driven Frameworks in Medical AI. Logic-based methods enhance transparency by decomposing tasks into interpretable modules (e.g., causal inference, treatment planning)(Bollikonda 2025; Rajpurkar et al. 2022). Yet, current systems often lack dynamic agent interactions and remain constrained by static task schemas. Bridging logical reasoning with adaptive multi-agent collaboration remains a core challenge for real-world clinical reasoning(Wu et al. 2024a; Su et al. 2024).
@@ 3.1 Problem Definition, Syllogism, and Logical Tree
In this work, each clinical QA task dataset $\mathcal{D}$ is defined as a set of tuples $\{(Q_{i},\mathcal{O}_{i},C_{i})\}_{i=1}^{N}$, where $Q_{i}=\text{`What is the correct diagnosis/management for the patient?'}$ is the question text, $\mathcal{O}_{i}=\{O_{i_{1}},O_{i_{2}},\dots,O_{i_{N_{K}}}\}$ is the set of candidate answers, $N_{K}$ is the number of candidate answers, and $C_{i}$ is the correct answer. Depending on the required answer, we consider Multiple-choice Question (MCQ), which means a candidate set $\{O_{i_{1}},O_{i_{2}},\dots,O_{i_{N_{K}}}\}$ is provided and the system must select the most appropriate option(s) (Pal, Umapathi, and Sankarasubbu 2022).
@@ 3.2 Agent Designs in MedLA
where $\mathcal{O}_{\mathrm{M^{(j)}\text{-}Agent}}(\cdot)$ is $j$-th M-Agent, and $\langle\texttt{pro}_{M}\rangle$ is the prompt template (in the Appendix). The prompt will guide the model to further split the subproblems and thus form a more complex logical tree $\mathcal{T}_{M^{(j)}}$. $\mathcal{T}_{M^{(j)}}$ contains a set of syllogism nodes $V=\{v_{1},v_{2},\dots,v_{i},\dots,v_{|N_{V}|}\}$ and a set of directed edges $E$. ${\mathcal{T}}_{\text{other}}$ is an optional input, which is used to provide the tree structure of other M-Agents in the previous round for agent-agent discussion. It is worth noting that in this paper, since LLMs cannot directly produce structured tree outputs, the $v_{i}$ we obtain by guiding LLM is a text block containing a syllogism(check Appendix C).
@@ 3.3 Logical Reasoning Workflow of MedLA
Method Reference MedDDx Benchmarks (Su et al. 2025a)
@@ 4. Experiments
Datasets & Benchmarks. We evaluate on three complementary benchmarks. (i) MedDDx benchmarks. To stress differential-diagnosis reasoning, we use MedDDx benchmark(Su et al. 2025a) : Tests differential diagnosis across Basic, Intermediate, and Expert tiers, with difficulty defined by the semantic similarity of distractors from STaRK-Prime (Wu et al. 2024b). (ii) Multi-choice medical QA benchmarks (Xiong et al. 2024). A suite combining MMLU-Med, MedQA-US, and BioASQ-Y/N to test general medical knowledge. It covers factual recall, guideline interpretation, and clinical decision-making. (iii) Expert-Level Medical Reasoning and Understanding benchmark(Zuo et al. 2025).To assess expert-level reasoning, we use MedXpertQA, a challenging and comprehensive benchmark for advanced medical knowledge (details in Table 6). The benchmark serves as our general-purpose test bed. The details of the benchmarks are shown in Appendix D.
Method Reference Multi-choice medical QA benchmarks (Xiong et al. 2024)
@@ Methods
Baselines. We benchmark MedLA against four representative paradigms. (i) Graph-based reasoning methods ground answers on biomedical knowledge graphs—QAGNN (Yasunaga et al. 2021), JointLK (Sun et al. 2022), and DRAGON (Yasunaga et al. 2022). (ii) Multi-agent voting systems aggregate LLM outputs without explicit logic trees—Majority Voting, DyLAN (Liu et al. 2024d), MedAgents (Tang et al. 2024), and MDAgents (Liu et al. 2024c). (iii) Stand-alone LLMs rely purely on parametric knowledge, including LLaMA-2-7B/13B, Mistral-7B, MedAlpaca-7B, PMC-LLaMA-7B, LLaMA 3-8B, LLaMA 3-OB-8B, and the stronger LLaMA 3.1-8B; we also report their chain-of-thought (CoT) variants. (iv) Retrieval-augmented generation (RAG) couples an LLM with external retrievers—Self-RAG (7B/13B) (Asai et al. 2023), KG-Rank (Yang et al. 2024), KG-RAG (Soman et al. 2023), and MedRAG (70B) (Xiong et al. 2024). Together, these baselines span parametric, retrieval-augmented, graph-grounded, and naïve multi-agent strategies, furnishing a comprehensive backdrop against which to gauge MedLA’s contributions. We did not include methods that require additional data and require fine-tuning of the larger model (e.g., KGAREVION (Su et al. 2025b)).
@@
Evaluation Metric & Testing Protocol & Implementation Details. Following (Su et al. 2025a), the performance is measured by accuracy (Acc), averaged over three independent runs ( $\pm$ std). For every run, we shuffle the benchmark order and reset the LLM’s sampling state. All experiments used the officially provided base model weights and configurations, and vLLM (v0.7.2). 8-card A100-80GB GPU servers are used for testing. The raw data of the dataset involved in the experiments is adopted from MIRAGE [^1] https://github.com/Teddy-XiongGZ/MIRAGE. More details can be found in the Appendix F.
@@
[Overall Performance Analysis] MedLA significantly outperforms baselines under open-source LLM settings. To evaluate the effectiveness of MedLA, we conducted comprehensive experiments on a series of medical reasoning benchmarks, all based on open-source large language models (e.g., LLaMA). To ensure the objectivity of the comparison, we directly refer to the baseline results reported in (Su et al. 2025a). The results are summarized in Table 1 and Table 2. We also provide a detailed summary in Table 9 and Table 10 in the Appendix. Analysis: (a) MedLA outperforms all baselines on both standard and challenging benchmarks, achieving state-of-the-art (SOTA) results across QA datasets and excelling in expert-level diagnosis on MedDDx. These results demonstrate that MedLA not only enhances factual reasoning but also improves differential diagnostic performance in real-world medical settings. (b) MedLA Beyond CoT and Baselines: MedLA surpasses both base and CoT-enhanced models under the same open-source foundation, showing that our logic extraction strategy effectively distills more reliable knowledge and supports dynamic reasoning. (c) MedLA Stronger than Multi-Agent Systems: MedLA outperforms multi-agent baselines, indicating that its logic tree enables deeper interaction and better coordination among agents even without explicit role-based decomposition. (d) MedLA Outperforms RAG without External Knowledge: MedLA exceeds RAG-based models despite not using external retrieval, proving its strong internal reasoning ability through structured logic alone.(e) MedLA Remains Effective Under New Challenges:On a newly introduced test set,MedXpertQA, MedLA’s performance remains outstanding, proving that its logic-enhanced reasoning ability is generalizable, rather than merely an optimization for known tasks, and is capable of effectively tackling new medical challenges.
@@
[Overall Performance Analysis] MedLA demonstrates strong performance advantages on commercial LLMs, validating its generality and robustness. Our previous experiments were primarily conducted on open-source large language models, where MedLA had already shown promising results across various medical reasoning tasks. To further assess the adaptability and competitiveness of MedLA under more powerful backbones, we conducted additional evaluations using state-of-the-art commercial LLMs such as DeepSeek. All baseline methods were also implemented on the same commercial model to ensure fair comparison. For objectivity, we referenced the baseline results reported in (Su et al. 2025a). The outcomes are presented in Table 3.
@@
Table 3: Performance comparison of MedLA and baseline on DeepSeek-based reasoning evaluated on the MedXpertQA benchmark(Zuo et al. 2025).
@@
[Detailed $N_{M}$ vs. Performance Analysis] MedLA leads to more effective collaboration between agents than the baseline multi-agent approach. To assess the contribution of our logic-tree protocol, we compare MedLA to a naïve ensemble that simply aggregates the same agents by majority vote on the MedDDx-Basic dataset with different LLM temperature parameters and number of agents. The LLaMA3.1-8B is used as the backbone for basic LLM systems. The results are shown in Fig. 4. Analysis: Across all configurations, MedLA yields higher accuracy and smaller error bars. With only seven agents($|N_{M}|{=}7$), it already surpasses the strongest majority voting baseline by $\approx 2$ pp; the gap widens to +6+6-77 pp at $|N_{M}|{=}17$. The default temperature for MedLA is 0.7, as we found that this parameter ensures both stability and sufficient room for discussion between different intelligences. Moreover, the narrower confidence intervals for MedLA confirm that premise-level arbitration not only elevates the mean performance but also stabilises it across random seeds.

### v2

@@ Abstract
Codehttps://github.com/alexander2618/MedLA
Extended versionhttps://arxiv.org/abs/2509.23725
@@ 1. Introduction
LLM has demonstrated significant advantages in the field of medical reasoning, and its deep learning architecture can efficiently extract knowledge from massive amounts of literature and clinical cases to provide intelligent assistance in diagnostic decision-making, and is expected to significantly improve the accessibility and popularization of medical knowledge(chang2024survey; kasneci2023chatgpt). Answering complex medical questions with large language models (LLMs) remains difficult. A practical system must integrate domain knowledge (liu2024ddk), patient information (yang2022large), and explicit logical reasoning (goh2024large). Recent general-purpose and domain-adapted models achieve strong open-benchmark scores (kim2024mdagents; tang2024medagents). In clinical use, however, they may hallucinate drug dosages, misapply guidelines, or draw invalid causal links, reducing diagnostic reliability (liu2023llavamed).
LLM-based medical reasoning follows two main paradigms. (a) Knowledge fine-tuning (singhal2023large; wang2025knowledge) retrains models on large medical corpora, improving accuracy at the cost of data, compute, and deployment agility. (b) Reasoning stimulation (lievin2024can) has been tried using multi-agent role-playing to accomplish tasks through discussion and cooperation, which is considered a flexible and low-cost solution (moor2023medflamingo; kim_mdagents_2024). However, we found that most current multi-agent systems only engage in positional discussions based on their rulings and cannot argue about logical details in depth. Current frameworks are not able to effectively localize logic/rule conflicts, thus be difficult to improve the performance.
Inspired by the ‘major premise-minor premise-conclusion’ paradigm of the classical syllogism, we use the syllogism as a minimal reasoning unit  (khemlani2012theories; jiang2023legal). Each triad consists of a generalized medical law (major premise), a patient-specific fact (minor premise), and the corresponding conclusion. By concatenating or parallelizing multiple triads, we construct an inference tree whose leaf nodes store empirical observations or domain rules, internal nodes hold intermediate inferences, and the root node gives the final clinical decision  (howson2005logic). Such inference trees offer two major advantages: (i) Traceability - each conclusion can be traced back to its supporting premises; (ii) Comparability - each intelligence can align the reasoning tree (logical tree) to pinpoint conflicts or omissions at the premise level. Embedding this explicit structure into a multi-agent framework provides an easily auditable logical template for systematic cross-intelligence error correction and consistency verification.
@@
The experimental results show that our MedLA outperforms existing MedDDx benchmarks (su2025kgarevion), medical QA benchmarks (xiong2024medrag), and medical reasoning benchmarks (zuo2025medxpertqa) by a large margin on both open-source and commercial LLM (Fig. 1(b)). Contributions. (a) We introduce MedLA, the first multi-agent framework for medical reasoning that represents each agent’s thought process as an explicit logical tree. This design enables fine-grained traceability of inferences and systematic detection of premise-level conflicts. (b) We develop a multi-round, graph-guided discussion mechanism in which agents iteratively compare and revise their logical trees, leading to robust cross-agent error correction and convergence to high-confidence, self-consistent reasoning structures. (c) We perform comprehensive evaluations on both differential diagnosis (MedDDx) and standard medical QA benchmarks, demonstrating that MedLA outperforms static role-based multi-agent systems and single LLM baselines.
@@ 3.1 Problem Definition, Syllogism, and Logical Tree
In this work, each clinical QA task dataset $\mathcal{D}$ is defined as a set of tuples $\{(Q_{i},\mathcal{O}_{i},C_{i})\}_{i=1}^{N}$, where $Q_{i}=\text{`What is the correct diagnosis/management for the patient?'}$ is the question text, $\mathcal{O}_{i}=\{O_{i_{1}},O_{i_{2}},\dots,O_{i_{N_{K}}}\}$ is the set of candidate answers, $N_{K}$ is the number of candidate answers, and $C_{i}\in\mathcal{O}_{i}$ is the ground-truth correct answer. The task for our system is to predict the correct answer from $\mathcal{O}_{i}$ given the input $(Q_{i},\mathcal{O}_{i})$. We primarily consider the Multiple-choice Question (MCQ) format (yun2023medmcqa).
@@ 3.2 Agent Designs in MedLA
where $\mathcal{O}_{\mathrm{M^{(j)}\text{-}Agent}}(\cdot)$ is $j$-th M-Agent, and $\langle\texttt{pro}_{M}\rangle$ is the prompt template (in the Appendix). The prompt will guide the model to further split the subproblems and thus form a more complex logical tree $\mathcal{T}_{M^{(j)}}$. $\mathcal{T}_{M^{(j)}}$ contains a set of syllogism nodes $V=\{v_{1},v_{2},\dots,v_{i},\dots,v_{|N_{V}|}\}$ and a set of directed edges $E$. ${\mathcal{T}}_{\text{other}}$ is an optional input, which is used to provide the tree structure of other M-Agents in the previous round for agent-agent discussion. It is worth noting that in this paper, since LLMs cannot directly produce structured tree outputs, the $v_{i}$ we obtain by guiding LLM is a text block containing a syllogism(check Appendix).
@@ 3.3 Logical Reasoning Workflow of MedLA
Method Reference MedDDx Benchmarks (su2025kgarevion)
@@ 4. Experiments
Datasets & Benchmarks. We evaluate on three complementary benchmarks. (i) MedDDx benchmarks. To stress differential-diagnosis reasoning, we use MedDDx benchmark(su2025kgarevion) : Tests differential diagnosis across Basic, Intermediate, and Expert tiers, with difficulty defined by the semantic similarity of distractors from STaRK-Prime (wu2024stark). (ii) Multi-choice medical QA benchmarks (xiong2024medrag). A suite combining MMLU-Med, MedQA-US, and BioASQ-Y/N to test general medical knowledge. It covers factual recall, guideline interpretation, and clinical decision-making. (iii) Expert-Level Medical Reasoning and Understanding benchmark(zuo2025medxpertqa).To assess expert-level reasoning, we use MedXpertQA, a challenging and comprehensive benchmark for advanced medical knowledge (details in Appendix). The benchmark serves as our general-purpose test bed. The details of the benchmarks are shown in Appendix.
Method Reference Multi-choice medical QA benchmarks (xiong2024medrag)
@@ Methods
Baselines. We benchmark MedLA against four representative paradigms. (i) Graph-based reasoning methods ground answers on biomedical knowledge graphs—QAGNN (yasunaga-etal-2021-qa), JointLK (sun-etal-2022-jointlk), and DRAGON (yasunaga2022dragon). (ii) Multi-agent voting systems aggregate LLM outputs without explicit logic trees—Majority Voting, DyLAN (liu2024dynamicllmpoweredagentnetwork), MedAgents (tang2024medagents), and MDAgents (liu2024dynamic). (iii) Stand-alone LLMs rely purely on parametric knowledge, including LLaMA-2-7B/13B, Mistral-7B, MedAlpaca-7B, PMC-LLaMA-7B, LLaMA 3-8B, LLaMA 3-OB-8B, and the stronger LLaMA 3.1-8B; we also report their chain-of-thought (CoT) variants. (iv) Retrieval-augmented generation (RAG) couples an LLM with external retrievers—Self-RAG (7B/13B) (asai2023selfrag), KG-Rank (yang2024kgrank), KG-RAG (soman2023kgrag), and MedRAG (70B) (xiong2024medrag). Together, these baselines span parametric, retrieval-augmented, graph-grounded, and naïve multi-agent strategies, furnishing a comprehensive backdrop against which to gauge MedLA’s contributions. We did not include methods that require additional data and require fine-tuning of the larger model (e.g., KGAREVION (su_kgarevion_2025)).
@@
Evaluation Metric & Testing Protocol & Implementation Details. Following (su2025kgarevion), the performance is measured by accuracy (Acc), averaged over three independent runs ( $\pm$ std). For every run, we shuffle the benchmark order and reset the LLM’s sampling state. All experiments used the officially provided base model weights and configurations, and vLLM (v0.7.2). 8-card A100-80GB GPU servers are used for testing. The raw data of the dataset involved in the experiments is adopted from MIRAGE [^1] https://github.com/Teddy-XiongGZ/MIRAGE. More details can be found in the Appendix.
@@
[Overall Performance Analysis] MedLA significantly outperforms baselines under open-source LLM settings. To evaluate the effectiveness of MedLA, we conducted comprehensive experiments on a series of medical reasoning benchmarks, all based on open-source large language models (e.g., LLaMA). To ensure the objectivity of the comparison, we directly refer to the baseline results reported in (su2025kgarevion). We also provide a detailed summary in the Appendix. Analysis: (a) MedLA outperforms all baselines on both standard and challenging benchmarks, achieving state-of-the-art (SOTA) results across QA datasets and excelling in expert-level diagnosis on MedDDx. These results demonstrate that MedLA not only enhances factual reasoning but also improves differential diagnostic performance in real-world medical settings. (b) MedLA Beyond CoT and Baselines: MedLA surpasses both base and CoT-enhanced models under the same open-source foundation, showing that our logic extraction strategy effectively distills more reliable knowledge and supports dynamic reasoning. (c) MedLA Stronger than Multi-Agent Systems: MedLA outperforms multi-agent baselines, indicating that its logic tree enables deeper interaction and better coordination among agents even without explicit role-based decomposition. (d) MedLA Outperforms RAG without External Knowledge: MedLA exceeds RAG-based models despite not using external retrieval, proving its strong internal reasoning ability through structured logic alone.(e) MedLA Remains Effective Under New Challenges:On a newly introduced test set,MedXpertQA, MedLA’s performance remains outstanding, proving that its logic-enhanced reasoning ability is generalizable, rather than merely an optimization for known tasks, and is capable of effectively tackling new medical challenges.
@@
[Overall Performance Analysis] MedLA demonstrates strong performance advantages on commercial LLMs, validating its generality and robustness. Our previous experiments were primarily conducted on open-source large language models, where MedLA had already shown promising results across various medical reasoning tasks. To further assess the adaptability and competitiveness of MedLA under more powerful backbones, we conducted additional evaluations using state-of-the-art commercial LLMs such as DeepSeek. All baseline methods were also implemented on the same commercial model to ensure fair comparison. For objectivity, we referenced the baseline results reported in (su2025kgarevion). The outcomes are presented in Table 3.
@@
Table 3: Performance comparison of MedLA and baseline on DeepSeek-based reasoning evaluated on the MedXpertQA benchmark(zuo2025medxpertqa).
@@
[Detailed $N_{M}$ vs. Performance Analysis] MedLA leads to more effective collaboration between agents than the baseline multi-agent approach. To assess the contribution of our logic-tree protocol, we compare MedLA to a naïve ensemble that simply aggregates the same agents by majority vote on the MedDDx-Basic dataset with different LLM temperature parameters and number of agents. The LLaMA3.1-8B is used as the backbone for basic LLM systems. The results are shown in Fig. 4. Analysis: Across all configurations, MedLA yields higher accuracy and smaller error bars. With only seven agents($|N_{M}|{=}7$), it already surpasses the strongest majority voting baseline by $\approx 2$ pp; the gap widens to +6+6-77 pp at $|N_{M}|{=}17$. The default temperature for MedLA is 0.7, as we found that this parameter ensures both stability and sufficient room for discussion between different intelligences.